OpenStudy (sissyedgar):

Which circles have an area of 16 Pi in2? Choose all answers that are correct. A. a circle with a radius of 2 in. B. a circle with a radius of 4 in. C. a circle with a diameter of 8 in. D. a circle with a diameter of 16 in.

OpenStudy (misty1212):

formula is \(\pi r^2\) so if the radius is \(2\) it would be \(\pi \times 2^2=4\pi\) so not that one !

OpenStudy (misty1212):

if the radius is \(4\) then it would be \(\pi\times 4^2=16\pi\) that is one of the right answers

OpenStudy (sissyedgar):

mhmm... is the other C?

OpenStudy (calculusxy):

Area = pi(r)^2

OpenStudy (sissyedgar):

...

OpenStudy (calculusxy):

Find the square root of 16.

OpenStudy (calculusxy):

That will be your radius.

OpenStudy (sissyedgar):

yah, 4

OpenStudy (calculusxy):

So it's B.

OpenStudy (calculusxy):

Great job!

OpenStudy (sissyedgar):

its more dan 1 answer!

OpenStudy (calculusxy):

I am not sure about the others :\

OpenStudy (sissyedgar):

i tink the other is C, its only 2!

OpenStudy (calculusxy):

Yeah I think you're correct

OpenStudy (calculusxy):

Half of 8 (diameter) is 4 (radius)

OpenStudy (radar):

The problem has only one correct answer. It doesn't say there is more than one correct answer. Don't read more into it, just select the correct answer.

OpenStudy (sissyedgar):

Choose All Answers That Are Correct! in my math, tht always means more than 1 answer!

OpenStudy (sissyedgar):

@StudyGurl14 2 sum up the chat, i tink its B & C

OpenStudy (studygurl14):

correct^

OpenStudy (sissyedgar):

O, YAY!

OpenStudy (sissyedgar):

4 more!

OpenStudy (radar):

Sorry, you are correct there are two correct answers and it is as you say B & C. I don't know where my head was for my post!